find the slope and y-intercept of the line which passes through the points (2,2) and (-1,-7). Then write the equation in slope intercept form.
Answer by rfer(16322)

slope=m
m=(-7-2)/(-1-2)
m=-9/-3
m=3
------------
y=mx+b
2=3*2+b
2=6+b
2-6=b
b=-4
-------------
y=3x-4